OpenStudy (anonymous):

Differentiate dy/dx = (x-4)(x^2 - 8x + 1)^1/2

OpenStudy (amorfide):

So you want d2y/dx2? use vdu+udv where V is the second term of the product, and U is the first term of the product. so U=x-4 V=(x^2-8x+1)^1/2

OpenStudy (anonymous):

Okay thank you. Sorry I have to go.